OpenStudy (anonymous):

Given the function f(x)=8 | x - 7 |. Calculate the following value: f(x+1)

OpenStudy (anonymous):

I know you just put x+1 where x is but i have no idea how to work with the absolute value f(x+1)=8|(x+1)-7|

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

U have done it right...f(x+1)=8|x-6|

OpenStudy (anonymous):

How did you get x-6? =p I feel kinda dumb right now..

OpenStudy (anonymous):

Nevermind...i was multiplying. Thank you

OpenStudy (anonymous):

ur welcome